OpenStudy (anonymous):

Solve 5√x−6=−20. WITH STEPS PLEASE IM CONFUSED. I know the answer is no solution, but why?

OpenStudy (anonymous):

first add 6 to both sides. u get 5√x=−14, then u divide 5 on both sides, u the get √x=−14/5, now the square root of a negative value (-14/5 is negative) doesn't exist

OpenStudy (anonymous):

the -6 is under the radical btw

OpenStudy (anonymous):

oh wut

OpenStudy (anonymous):

what then there's a solution...

OpenStudy (anonymous):

like it's \[5\sqrt{x-6}=-20\]

OpenStudy (anonymous):

5√x−6=−20 (try to isolate for x so take 5 to other side by dividing it on both sides) √x−6=−20/5 √x−6=-4 (√x−6)^2=(−20)^2 (square both sides to get rid of square root) x-6=400 x=406 (there is a solution)

OpenStudy (anonymous):

I got 22 the first time, but apparently that's wrong

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

I am so confused. I mean i guess it could be, but it wasn't an option. The right answer was apparently no solution

OpenStudy (anonymous):

are you sure? doesn't make sense...that's the way i would do it

OpenStudy (anonymous):

I don't understand how you did step three in your solution. why is it -20 on the right?

OpenStudy (anonymous):

think he messed up...

OpenStudy (anonymous):

Because that's the way that you stated the equation, with -20 on the right.. :P

OpenStudy (anonymous):

Ok thanks, I guess he just messed it up. THANK YOU GUYS!!

OpenStudy (anonymous):

divide by 5, square both sides, then solve for x

OpenStudy (anonymous):

x = 10

OpenStudy (anonymous):

@monkey220ms but you divided it by 5 to take the 5 off of the left side. which makes it -4

OpenStudy (anonymous):

@11calcBC wouldn't you get 22??

OpenStudy (anonymous):

Oh I see what your saying 5√x−6=−20 √x−6=−20/5 √x−6=-4 (√x−6)^2=(4)^2 x-6=16 x=22 This is what @rimshaa did but with the right numbers:) so it is 22

OpenStudy (anonymous):

Ok thanks! I'm gonna close it now.

OpenStudy (anonymous):

oh lol my bad and yes @monkey220ms has the right answer

OpenStudy (anonymous):

lol mixed up the + and -